two concentric circular coils of ten turns each are situated in the same plane.their radii are 20cm and 40cm and they carry respectively 0.2 and 0.3 ampere current in opposite direction.the magnitude in weber/m2 at the center in terms of permiabitity in vaccum is
1]35/4
2]1/80
3]7/80
4]5/4
 

Grade:8

2 Answers

Vikas TU
6 years ago
Radius of first coil, R1 = 20cm
            Radius of second coil, R2 = 40cm
            Number of turns = 10
            Current in first coil, I1 = 0.2 A
            Current in first coil, I2 = 0.3 A
 NI/2Ro            The formula of magnetic field at the center is μ            As the direction of current is opposite in both of the coils, they will generate opposite magnetic field.            So,                         B = B1 - B2)2/R2 – I1/R1N/2)(Io                        = (μ10/2)(0.2/0.2 – 0.3/0.4)o                        = (μ/4o                        = 5μ            So, the answer is option #4
Vikas TU
6 years ago
Radius of first coil, R1 = 20cm
            Radius of second coil, R2 = 40cm
            Number of turns = 10
            Current in first coil, I1 = 0.2 A
            Current in first coil, I2 = 0.3 A
 NI/2R. As the direction of current is opposite in both of the coils, they will generate opposite magnetic field.o     The formula of magnetic field at the center is μ            So,             B = B1 - B2)2/R2 – I1/R1N/2)(Io         = (μ10/2)(0.2/0.2 – 0.3/0.4)o          = (μ/4o         = 5μ      So, the answer is option #4